OpenStudy (anonymous):

Through (6, 7); parallel to the line passing through (7, 5) and (3, 1)

OpenStudy (anonymous):

m=(1-5)/(3-7)=-4/-4=1 Y-7=1(X-6) SIMPLIFY.

OpenStudy (anonymous):

thank you but i did that when i put in x-y+6=0 it say its wrong

OpenStudy (anonymous):

y-7=1(x-6) y-7=x-6 +7 +7 y=x+1

OpenStudy (anonymous):

thank you